# Towards agent-based-model informed neural networks

### v1

@@ Abstract
In this article, we present a framework for designing neural networks that remain consistent with the underlying principles of agent-based models. We begin by highlighting the limitations of standard neural differential equations in modeling complex systems, where physical invariants (like energy) are often absent but other constraints (like mass conservation, network locality, bounded rationality) must be enforced. To address this, we introduce Agent-Based-Model informed Neural Networks (ABM-NNs), which leverage restricted graph neural networks and hierarchical decomposition to learn interpretable, structure-preserving dynamics. We validate the framework across three case studies of increasing complexity: (i) a generalized Generalized Lotka–Volterra system, where we recover ground-truth parameters from short trajectories in presence of interventions; (ii) a graph-based SIR contagion model, where our method outperforms state-of-the-art graph learning baselines (GCN, GraphSAGE, Graph Transformer) in out-of-sample forecasting and noise robustness; and (iii) a real-world macroeconomic model of the ten largest economies, where we learn coupled GDP dynamics from empirical data and demonstrate gradient-based counterfactual analysis for policy interventions.
@@ Introduction
ensuring energy conservation and structure-preserving predictions even over long time horizons. This inductive bias distinguishes HNNs from standard neural ODE frameworks, offering improved generalization and interpretability in modeling physical systems. See Fig. 1 right, which shows that Hamiltonian neural networks can preserve energy conservation, and thus physical realism. Analogously, one can encode physical reality via Lagrangian Neural Networks that parameterize the Lagrangian and derive dynamics from the Euler–Lagrange equations [Cranmer2020LNN]. Related physics-informed approaches constrain neural surrogates to satisfy governing equations, either through PDE residuals in physics-informed neural networks [Raissi2019PINN] or through operator-learning architectures such as Fourier Neural Operators [Li2020FNO], and have been surveyed under the broader banner of physics-informed machine learning [Karniadakis2021PIML]. by using pre-defined dictionary of functions.
@@
However, many socio-techno-economic systems operate far from equilibrium, and this is particularly evident in systems exhibiting self-organized criticality, where the endogenous drive for efficiency pushes the system to a fragile state at the edge of chaos, making large-scale fluctuations an intrinsic feature rather than an anomaly [bouchaud2024self]. To model the dynamics of socio-techno-economic systems that operate far from equilibrium, a diverse array of non-equilibrium frameworks has been developed, ranging from (i) stochastic frameworks, (ii) computational & simulation frameworks, to (iii) macroscopic deterministic approximations. Stochastic frameworks provide the exact mathematical description of a system’s evolution by tracking probability distributions, event timings, or network states without ignoring fluctuations. Central to this approach is the Master Equation, which describes the time-evolution of the probability of the system being in a specific discrete microstate and serves as the fundamental formalism for stochastic jump processes in continuous time [vankampen1992stochastic, gardiner2009stochastic]. Complementing this state-based view are Self-Exciting Point Processes, such as Hawkes Processes, which model the precise timing of events in continuous time. These capture how past events trigger “avalanches” of future activity, a feature crucial for modeling endogenous bursting in financial and social systems [hawkes1971spectra, bacry2015hawkes]. Additionally, for systems defined by sparse, heterogeneous networks, analytical techniques like Cavity Methods or Message Passing (e.g., Belief Propagation) provide tools to solve statistical mechanics problems by capturing how local topology influences marginal probabilities [mezard2009information, pearl1988probabilistic]. Since exact solution to the Master Equation are often analytically insoluble for large complex systems, computational and simulation frameworks are employed to simulate specific realizations of system dynamics. The primary tool for capturing bounded rationality, heterogeneity, and complex emergence from the bottom up is Agent-Based Modeling (ABM), which simulates the granular, adaptive behavior of autonomous entities [epstein1996growing, farmer2024quantitative]. For a more rigorous physical approach, Kinetic Monte Carlo (KMC) algorithms solve the Master Equation numerically by generating exact stochastic trajectories via event-by-event sampling based on physical rates [gillespie1976general, voter2007introduction]. Furthermore, Directed Percolation provides a geometric simulation approach to identify critical tipping points and connectivity thresholds in non-equilibrium spreading processes, distinguishing between local containment and systemic contagion [stauffer2018introduction, hinrichsen2000non]. Finally, deterministic and aggregate frameworks offer macroscopic approximations by simplifying the system to ignore fluctuations and correlations, describing dynamics in the limit of infinite system size. Evolutionary Population Dynamics, encompassing Generalized Lotka-Volterra equations and Replicator Dynamics, model the non-linear competition, selection, and cyclic dominance between aggregate populations driven by fitness differences [hofbauer1998evolutionary, nowak2006evolutionary]. Similarly, Mean-Field ODE Approximations are derived by averaging the Master Equation to describe the time-evolution of system densities, assuming that every agent interacts with the certain averaged variables [barrat2008dynamical, pastorsatorras2001epidemic].
@@
In this paper, we will introduce abm-inspired neural networks that use the neural ODE framework with inductive biases from agent-based models to model out-of-equilibrium dynamics of complex systems. Central inductive bias we use is the sepration of laws into self-dynamics neural network and restricted graph neural network. Graph neural networks (GNNs) make these relational inductive biases explicit and have been used as learnable physics engines for interaction-driven systems [Battaglia2018Relational, SanchezGonzalez2019GraphPhysics]. Restricted graph neural networks are a special class of graph neural networks that are restricted to only use the immediate neighbors of a node to compute the output [Vasiliauskaite2024Generalization]. Graph Neural Networks operate as learnable message passing systems, where the exchange of information between nodes/agents is governed by differentiable neural functions rather than fixed analytical equations [gilmer2017neural]. Thus, a GNN enables the system to “learn physics”: instead of relying on a human derivation of how agents should influence each other, the network explicitly learns the non-linear interaction rules that best predict the system’s evolution directly from data. Finally, by analogy to HNNs for Neural ODEs, we introduce additional structure to neural differential equations to model different conservation laws of agent-based models.
In summary, this work: (i) introduces ABM-informed Neural Networks (ABM-NNs) for modeling complex systems while preserving structural constraints; (ii) validates the approach on a generalized GLV system, recovering analytical laws from data; (iii) demonstrates superior out-of-sample generalization compared to GCN, SAGE, and GraphGPS baselines on contagion dynamics; and (iv) applies the framework to empirical macroeconomic data, enabling counterfactual policy analysis via gradient-based interventions.
@@ Methods
Note that a restricted graph neural network is needed [Vasiliauskaite2024Generalization] to enforce the correct inductive biases for the vector field: within infinitesimal time, a node can interact only with its immediate neighbors. A restricted graph neural network is a graph neural network that is restricted to only use the immediate neighbors of a node to compute the output (i.e. local interaction function). In this paper, we use the terms restricted graph neural network and Graph Neural Vector Field interchangeably. We will only use the local interaction functions, and omit the word local for simplicity. The Graph Neural Vector Field framework can better capture the vector field of dynamics out of sample than a pure feed-forward neural network or even state-of-the-art graph neural networks [Vasiliauskaite2024Generalization], because it incorporates an appropriate set of inductive biases. However, how to proceed in the case of general dynamics with multiple states is not clear. By analogy with Hamiltonian Neural Networks for Neural ODEs, we propose agent-based-model informed neural networks (ABM-NNs) that model the dynamics of complex systems, by inorporating principles of agent-based models into the neural networks. These principles can be connected to the laws of conservation of mass, interaction between agents, or other principles of agent-based models.
@@
Learning fuctionals $F$, $G$, $H$ can be implemented in differnt ways. By analogy to Fourier, Kernel methods, we can represent $F$, $G$, $H$ as linear combinations of the learned functions $\phi_{i}$ and $\psi_{i}$:
@@
, where $\langle\cdot\rangle$ denotes the dot-product $\langle\textbf{c}_{f}\phi\rangle=\sum_{i=1}^{k}c_{f,i}\cdot\phi_{i}$. Similarly, we can represent $G=\langle\textbf{c}_{g}\phi\rangle+\langle\textbf{c}_{g}\psi\rangle$, and $H=\langle\textbf{c}_{h}\phi\rangle+\langle\textbf{c}_{h}\psi\rangle$. Finally, we would like to stress, that different case-studies require slightly different inductive-biases in order to better capture agent-based behavour in different complex systems, see results sections for more details.
@@ Results
In the results section we showcase the proposed framework through three case studies, each relaxing more assumptions than the previous one. The first study revisits a generalized Lotka–Volterra system with analytical dynamics, where only a handful of parameters are calibrated. It adapts population dynamics to a fictional cosmological scenario featuring five factions competing for resources, with $F=\phi+\psi$ fixed analytically and finite resources or carrying capacity is enforced via a predefined adjacency matrix with non-zero diagonals. The second study focuses on an SIR contagion model on graphs. We begin with analytical $F,G,H$ and learned neural operators $\phi,\psi$, then benchmark against GCN [Kipf2017GCN], SAGE [Hamilton2017GraphSAGE], and GraphGPS Transformer [Rampasek2022RecipeGPS] baselines, and additionally test robustness to observational noise, new initial conditions,, evaluate out-of-sample generalization on unseen graphs, and finally learn the linear functionals directly from data to handle interventions. Finall third study targets latent macroeconomic dynamics of top 10 largest economies in the world, where empirical time series inform both the interaction matrix and the subsequent counterfactual simulations of the learned system. In the macroeconomic model, we put special emphasis on heterogeneity by using Feature-wise Linear Modulation adapters (FiLM).
It is important to note that the framework presented in the methods section incorporates several explicit and implicit inductive biases relevant to the design of agent-based models. We establish spatiotemporal constraints by using restricted Graph Neural Networks to prioritize local peer interactions over global knowledge, while leveraging the discretization step of Neural ODEs to model continuous decision-making and strictly define the speed of information processing. To capture human-like limitations, we utilize gradient descent to naturally seek satisficing heuristics rather than global optima. We combine this computational constraint with FiLM-based heterogeneity, effectively replacing the uniform representative agent with a diverse population. These behavioral mechanisms are further grounded by strict conservation laws and an empirical prior, ensuring that the agents’ evolution remains physically consistent and aligned with historical data rather than abstract theory. Nevertheless, we also demonstrate that the framework remains sufficiently expressive to capture complex dynamics in scenarios where we simulate ground-truth behavior.
@@
We first take a well-studied analytical GLV model and transplant it into the fictional three-body problem trilogy [Liu2008ThreeBody], where distinct sociopolitical factions (the mainstream Planetary Defense Council, the militaristic Solar Fleet, the strategic Wallfacers, the traitorous ETO, and the opportunistic Escapists) compete for limited industrial capacity and scientific resources to either save or destroy humanity. Instantiating the GLV template from Eqs. (3)–(10) in this narrative gives the most constrained experiment in our suite: every term in the right-hand side is analytical, so the auto-diff framework [Paszke2019PyTorch] only calibrates the intrinsic growth parameters $r_{i}$ while holding the pairwise law $\phi_{\text{pair}}(x_{i},x_{j})=x_{i}x_{j}$ fixed. Each faction $i$ is treated as a node with capacity $X_{i}(t)$, restricted potential $r_{i}$, and interaction weights $a_{ij}$, yielding the additive self-plus-interaction decomposition. Data generation process is defined as following:
@@
where $S_{i}(t)$ encodes Sophon-imposed physics access (extraterrestrial supercomputers that lock down human scientific progress in the novel), $\tau_{i}(t)$ is an exogenous policy regime, and $a_{ij}$ captures sabotage, cooperation, or spillovers. Crucially, non-zero diagonal elements $a_{ii}\neq 0$ imply finite resources, effectively enforcing a carrying capacity that prevents unbounded growth. In the first case study, we want to show how ABM-inspired mean-field differential equations can be calibrated, even if different exogenous interventions are being applied. Optimization is performed with gradient descent-type optimizer, over the computational graph, that is unfolded with system of oridnary differential equations.
@@
where $\phi_{i}(x_{i})=r_{i}x_{i}$ and $\psi(\textbf{x},A)=\sum_{j=1}^{N}a_{ij}x_{i}x_{j}$. We can think of $\phi_{i}{\theta}$ as very simple differentiable network that has a single parameter $r_{i}$ and $\psi_{\theta}$ as the interaction function with no-free parameters. We start with this simple case, since it is very close to classical ABM model. But the calibration of parameters, will be done by using automatic differentiation framework, over the differentiable ODE solver.
In this fictional story, different fractions are competing for resources, and different exogenous effects are happening. One is through the use of multiplicative factor $S_{i}(t)$, and second one via the use of additive factor $\tau_{(}t)$. Both of this factor are affecting the effective growth rate of different fractions in the generalized Lotka–Volterra system (GLV). We will now describe exogenous effects, that we will use in our GLV system.
@@
Regularization weights $\lambda_{\mathrm{axis}}=1$, $\lambda_{\mathrm{origin}}=1$ with a warmup of 100 epochs. In Appendix Fig. 5, we show the training results for the ABM-informed neural network on SIR dynamics. In this experiment (Appendix Fig. 5), we generate a single Erdős–Rényi contact network with node count $n=100$ and edge probability $p=0.05$. Ground-truth macroscopic curves $s(t)=\sum_{i}S_{i}(t)$, $i(t)=\sum_{i}I_{i}(t)$, and $r(t)=\sum_{i}R_{i}(t)$ are obtained by simulating an SIR model with infection and recovery rates $\beta=0.4$ and $\gamma=0.2$, respectively, using RK4 time stepping with $\mathrm{d}t=0.1$ up to $t_{\max}=60$. Initial conditions are set by infecting up to 50 nodes sampled from the largest connected component, with the remainder susceptible and no recoveries: $S_{i}(0)\in\{0,1\}$, $I_{i}(0)\in\{0,1\}$, $R_{i}(0)=0$, and $S_{i}(0)+I_{i}(0)=1$. Training uses a differentiable RK4 rollout in PyTorch on the fixed graph and minimizes an integral L1 path loss between the stacked state trajectories with a simple curriculum on the rollout horizon: starting at $t=1$ and increasing by $\Delta t=1$ every 50 epochs up to a cap (default $t_{\text{train}}\leq 30$). We add a regularization functional $\Omega(\phi_{1},\phi_{2})$ and optimize the sum of the data loss and $\Omega$. Appendix Fig. 5 also shows the training losses and regularizers at different epochs. We observe that the model learns the dynamics of the system, and the regularization terms enforce the desired properties. Other experimental settings are detailed in Appendix Appendix B: Contagion dynamics – case study.
After training, we simulate the learned ABM-informed neural network over the full horizon to compare $\tilde{s}(t),\tilde{i}(t),\tilde{r}(t)$ with the ground truth, inspect 1D slices of $\phi_{1}(0.5,x)$. Importantly, the structure of the SIR equations preserves node-wise mass: for each $i$, $\dot{S}_{i}+\dot{I}_{i}+\dot{R}_{i}=0$. With simplex-consistent initial conditions $S_{i}(0)+I_{i}(0)+R_{i}(0)=1$, the conservation law implies $S_{i}(t)+I_{i}(t)+R_{i}(t)=1$ for all $t$. Our architectural choices (nonnegative fluxes, zero-at-zero behavior) support this invariance, and the numerical scheme maintains it through per-node renormalization during simulation. In Hamiltonian neural networks, the neural parameterization enforces a symplectic vector field, which preserves a first integral (the Hamiltonian/energy) by construction. Analogously, in our epidemic setting the compartmental structure induces a conservation law (mass on each node). By designing the neural right-hand side to respect the compartmental flux decomposition (nonnegative edge transmissions and nonnegative recoveries with zero flux at zero state), we embed an inductive bias that is the mass-conserving counterpart of the energy-preserving bias in HNNs. Thus, while HNNs preserve energy through symplectic structure, our nonnegative flux parameterizations preserve mass through a discrete continuity constraint, ensuring that neural dynamics remain on the probability simplex for each agent.
@@
All models are trained for 1000 epochs using the Adam optimizer with a learning rate of $10^{-4}$, minimizing the macroscopic $L_{1}$ loss derived from a differentiable RK4 rollout. ABM-NN Architecture: The ABM-NN decomposes dynamics into pairwise interactions ($\phi_{1}$) and self-recovery ($\phi_{2}$), both implemented as 3-layer MLPs (hidden dim: 32) with LeakyReLU activations. Baseline Architectures: The GCN baseline employs a stack of two GCNConv layers (hidden dimension 64) followed by a linear head. Similarly, GraphSAGE uses two SAGEConv layers (mean aggregation, hidden dimension 64) with a linear head. The GraphGPS transformer processes inputs through two GPS layers—each combining a local GCNConv with a 4-head global attention mechanism (hidden dimension 64)—before a final projection head. All baselines output a 3D time derivative vector for the RK4 solver.
@@
$\dot{\textbf{x}}=\underbrace{\phi_{1}(x_{i},y_{i},e_{i})}_{\text{self-interaction}}+\underbrace{\psi_{1}(\textbf{x},\textbf{A},e_{i},e_{j})}_{\text{interaction term}},$ (30)
$\dot{\textbf{y}}=\underbrace{\phi_{3}(y_{i},x_{i},e_{i})}_{\text{self-interaction}},$ (31)
@@
$\psi_{1}(x_{i},\textbf{x},\textbf{A},e_{i},e_{j})=\sum_{j}A_{ij}\,\phi_{2}(x_{i},x_{j},e_{i},e_{j}),$ (32)
where $\phi_{2}$ is a feed-forward neural network, and A is the adjacency matrix of the graph.
@@
The shared self map $\phi_{1}$ follows the decomposition of Eq. (33),
@@
where $\phi_{g}(\cdot)$ is a feed-forward neural network that models annualized growth and $\phi_{d}(\cdot)$ is a feed-forward neural network that captures structural drifts. The latent macro network $\phi_{3}$ is a feed-forward neural network that models the dynamics of the macro channels. For economic modeling and macroeconomic data scarcity, we would like to have parsimonious models, we will use a shared feed-forward neural network for both $\phi_{g}(\cdot)$ and $\phi_{d}(\cdot)$ and $\phi_{3}$. Shared models correspond to the idea that economic laws follow some hidden economic universality laws or principles, and the same economic laws apply to all countries. However, in reality, each country has additional hidden variables or factors that are not captured by the shared model, and thus we will use FiLM-style adapters to model the country-specific conditional dynamics. Feature-wise Linear Modulation adapters (FiLM) were originally proposed for visual reasoning [perez2018film], and they have been shown to be effective in learning conditional models. In our case, we will use FiLM-style adapters to model the country-specific conditional macro dynamics.

### v2

@@ Abstract
In this article, we present a framework for designing neural networks that remain consistent with the underlying principles of agent-based models. We begin by highlighting the limitations of standard neural differential equations in modeling complex systems, where physical invariants (like energy) are often absent but other constraints (like mass conservation, information locality, bounded rationality) must be enforced. To address this, we introduce Agent-Based-Model informed Neural Networks (ABM-NNs), which leverage restricted graph neural networks and hierarchical decomposition to learn interpretable, structure-preserving dynamics. We validate the framework across three case studies of increasing complexity: (i) a Generalized Lotka–Volterra system, where we recover ground-truth parameters from short trajectories in presence of interventions; (ii) a graph-based SIR contagion model, where our method outperforms state-of-the-art graph learning baselines (GCN, GraphSAGE, Graph Transformer) in out-of-sample forecasting and noise robustness; and (iii) a real-world macroeconomic model of the ten largest economies, where we learn coupled GDP dynamics from empirical data and demonstrate counterfactual analysis for policy interventions.
@@ Introduction
ensuring energy conservation and structure-preserving predictions even over long time horizons. This inductive bias distinguishes HNNs from standard neural ODE frameworks, offering improved generalization and interpretability in modeling physical systems. See Fig. 1 right, which shows that Hamiltonian neural networks can preserve energy conservation, and thus physical realism. Analogously, one can encode physical reality via Lagrangian Neural Networks that parameterize the Lagrangian and derive dynamics from the Euler–Lagrange equations [Cranmer2020LNN]. Related physics-informed approaches constrain neural surrogates to satisfy governing equations, either through PDE residuals in physics-informed neural networks [Raissi2019PINN] or through operator-learning architectures such as Fourier Neural Operators [Li2020FNO], and have been surveyed under the broader banner of physics-informed machine learning [Karniadakis2021PIML].
@@
However, many socio-techno-economic systems operate far from equilibrium, and this is particularly evident in systems exhibiting self-organized criticality, where the endogenous drive for efficiency pushes the system to a fragile state at the edge of chaos, making large-scale fluctuations an intrinsic feature rather than an anomaly [bouchaud2024self]. To model the dynamics of socio-techno-economic systems that operate far from equilibrium, a diverse array of non-equilibrium frameworks has been developed, ranging from (i) stochastic frameworks, (ii) computational & simulation frameworks, to (iii) macroscopic deterministic approximations. Stochastic frameworks provide the exact mathematical description of a system’s evolution by tracking probability distributions, event timings, or network states without ignoring fluctuations. Central to this approach is the Master Equation, which describes the time-evolution of the probability of the system being in a specific discrete microstate and serves as the fundamental formalism for stochastic jump processes in continuous time [vankampen1992stochastic, gardiner2009stochastic]. Complementing this state-based view are Self-Exciting Point Processes, such as Hawkes Processes, which model the precise timing of events in continuous time. These capture how past events trigger “avalanches” of future activity, a feature crucial for modeling endogenous bursting in financial and social systems [hawkes1971spectra, bacry2015hawkes]. Additionally, for systems defined by sparse, heterogeneous networks, analytical techniques like Cavity Methods or Message Passing (e.g., Belief Propagation) provide tools to solve statistical mechanics problems by capturing how local topology influences marginal probabilities [mezard2009information, zdeborova2016statistical]. Since exact solution to the Master Equation are often analytically insoluble for large complex systems, computational and simulation frameworks are employed to simulate specific realizations of system dynamics. The primary tool for capturing bounded rationality, heterogeneity, and complex emergence from the bottom up is Agent-Based Modeling (ABM), which simulates the granular, adaptive behavior of autonomous entities [epstein1996growing, farmer2024quantitative]. For a more rigorous physical approach, Kinetic Monte Carlo (KMC) algorithms solve the Master Equation numerically by generating exact stochastic trajectories via event-by-event sampling based on physical rates [gillespie1976general, voter2007introduction, bottcher2020unifying]. Furthermore, Directed Percolation provides a geometric simulation approach to identify critical tipping points and connectivity thresholds in non-equilibrium spreading processes, distinguishing between local containment and systemic contagion [stauffer2018introduction, hinrichsen2000non]. Finally, deterministic and aggregate frameworks offer macroscopic approximations by simplifying the system to ignore fluctuations and correlations, describing dynamics in the limit of infinite system size. Evolutionary Population Dynamics, encompassing Generalized Lotka-Volterra equations and Replicator Dynamics, model the non-linear competition, selection, and cyclic dominance between aggregate populations driven by fitness differences [hofbauer1998evolutionary, nowak2006evolutionary]. Similarly, Mean-Field ODE Approximations are derived by averaging the Master Equation to describe the time-evolution of system densities, assuming that every agent interacts with the certain averaged variables [barrat2008dynamical, pastorsatorras2001epidemic].
@@
In this paper, we will introduce abm-inspired neural networks that use the neural ODE framework with inductive biases from agent-based models to model out-of-equilibrium dynamics of complex systems. Central inductive bias we use is the separation of laws into self-dynamics neural network and restricted graph neural network. Graph neural networks (GNNs) make these relational inductive biases explicit and have been used as learnable physics engines for interaction-driven systems [Battaglia2018Relational, SanchezGonzalez2019GraphPhysics]. Restricted graph neural networks are a special class of graph neural networks that are restricted to only use the immediate neighbors of a node to compute the output [Vasiliauskaite2024Generalization]. Graph Neural Networks operate as learnable message passing systems, where the exchange of information between nodes/agents is governed by differentiable neural functions rather than fixed analytical equations [gilmer2017neural]. Thus, a GNN enables the system to “learn physics”: instead of relying on a human derivation of how agents should influence each other, the network explicitly learns the non-linear interaction rules that best predict the system’s evolution directly from data. Finally, by analogy to HNNs for Neural ODEs, we introduce additional structure to neural differential equations to model different conservation laws of agent-based models.
In summary, this work: (i) introduces ABM-informed Neural Networks (ABM-NNs) for modeling complex systems while preserving structural constraints; (ii) validates the approach on a generalized GLV system, recovering exact parameters of analytical laws from simulated data; (iii) demonstrates superior out-of-sample generalization compared to GCN, SAGE, and GraphGPS baselines on contagion dynamics; and (iv) applies the framework to empirical macroeconomic data, enabling counterfactual policy analysis via simulation.
@@ Methods
Note that a restricted graph neural network is needed [Vasiliauskaite2024Generalization] to enforce the correct inductive biases for the vector field: within infinitesimal time, a node can interact only with its immediate neighbors. A restricted graph neural network is a graph neural network that is restricted to only use the immediate neighbors of a node to compute the output (i.e. local interaction function). In this paper, we use the terms restricted graph neural network and Graph Neural Vector Field interchangeably. The Graph Neural Vector Field framework can better capture the vector field of dynamics out of sample than a pure feed-forward neural network or even state-of-the-art graph neural networks [Vasiliauskaite2024Generalization], because it incorporates an appropriate set of inductive biases. However, how to proceed in the case of general dynamics with multiple states is not clear. By analogy with Hamiltonian Neural Networks for Neural ODEs, we propose agent-based-model informed neural networks (ABM-NNs) that model the dynamics of complex systems, by incorporating principles of agent-based models into the neural networks. These principles can be connected to the laws of conservation of mass, interaction between agents, or other principles of agent-based models.
@@
Learning functionals $F$, $G$, $H$ can be implemented in different ways. By analogy to Fourier, Kernel methods, we can represent $F$, $G$, $H$ as linear combinations of the learned functions $\phi_{i}$ and $\psi_{i}$:
@@
where $\langle\cdot\rangle$ denotes the dot-product $\langle\textbf{c}_{f}\phi\rangle=\sum_{i=1}^{k}c_{f,i}\cdot\phi_{i}$. Similarly, we can represent $G=\langle\textbf{c}_{g}\phi\rangle+\langle\textbf{c}_{g}\psi\rangle$, and $H=\langle\textbf{c}_{h}\phi\rangle+\langle\textbf{c}_{h}\psi\rangle$. Finally, we would like to stress, that different case-studies require slightly different inductive-biases in order to better capture agent-based behavour in different complex systems, see results sections for more details.
@@ Results
In the results section we showcase the proposed framework through three case studies, each relaxing more assumptions than the previous one. The first study revisits a generalized Lotka–Volterra system with analytical dynamics, where only a handful of parameters are calibrated. It adapts population dynamics to a fictional cosmological scenario featuring five factions competing for resources, with $F=\phi+\psi$ fixed analytically and finite resources or carrying capacity is enforced via a predefined adjacency matrix with non-zero diagonals. The second study focuses on an SIR contagion model on graphs. We begin with analytical $F,G,H$ and learned neural operators $\phi,\psi$, then benchmark against GCN [Kipf2017GCN], SAGE [Hamilton2017GraphSAGE], and GraphGPS Transformer [Rampasek2022RecipeGPS] baselines, and additionally test robustness to observational noise, new initial conditions, evaluate out-of-sample generalization on unseen graphs, and finally learn the linear functionals directly from data. Finally, third study targets latent macroeconomic dynamics of top 10 largest economies in the world, where empirical time series inform both the interaction matrix and the subsequent counterfactual simulations of the learned system. In the macroeconomic model, we put special emphasis on heterogeneity by using Feature-wise Linear Modulation adapters (FiLM) [perez2018film].
It is important to note that the framework presented in the methods section incorporates several explicit and implicit inductive biases relevant to the design of agent-based models. We establish spatiotemporal constraints by using restricted Graph Neural Networks to prioritize local peer interactions over global knowledge, while leveraging the discretization step of Neural ODEs to model the speed of information processing. To capture human-like limitations, we utilize gradient descent to naturally seek satisficing heuristics rather than global optima. We combine this computational constraint with FiLM-based heterogeneity, effectively replacing the uniform representative agent with a diverse population. These behavioral mechanisms are further grounded by strict conservation laws and an empirical prior, ensuring that the agents’ evolution remains physically consistent and aligned with historical data rather than abstract theory. Nevertheless, we also demonstrate that the framework remains sufficiently expressive to capture complex dynamics in scenarios where we simulate ground-truth behavior.
@@
We first take a well-studied analytical GLV model and apply it into the fictional three-body problem trilogy [Liu2008ThreeBody], where distinct sociopolitical factions (the mainstream Planetary Defense Council, the militaristic Solar Fleet, the strategic Wallfacers, the traitorous ETO, and the opportunistic Escapists) compete for limited industrial capacity and scientific resources to either save or destroy humanity. Each faction $i$ is treated as a node with capacity $X_{i}(t)$, restricted potential $r_{i}$, and interaction weights $a_{ij}$, yielding the additive self-plus-interaction decomposition. Data generation process is defined as following:
@@
where $S_{i}(t)$ encodes Sophon-imposed physics access (extraterrestrial supercomputers that lock down human scientific progress in the novel), $\tau_{i}(t)$ is an exogenous policy regime, and $a_{ij}$ captures sabotage, cooperation, or spillovers. Crucially, non-zero diagonal elements $a_{ii}\neq 0$ imply finite resources, effectively enforcing a carrying capacity that prevents unbounded growth. In the first case study, we want to show how ABM-inspired differential equations can be calibrated, even if different exogenous interventions are being applied. Optimization is performed with gradient descent-type optimizer, over the computational graph, that is unfolded with system of ordinary differential equations.
@@
where $\phi_{1}(x_{i})=r_{i}x_{i}$ and $\psi(\textbf{x},A)=\sum_{j=1}^{N}a_{ij}x_{i}x_{j}$. We can think of $\phi_{1}$ as very simple differentiable network that has a single parameter $r_{i}$ and $\psi$ as the interaction function with no-free parameters. We start with this simple case, since it is very close to classical ABM model. But the calibration of parameters, will be done by using automatic differentiation framework, over the differentiable ODE solver.
In this fictional story, different factions are competing for resources, and different exogenous effects are happening. One is through the use of multiplicative factor $S_{i}(t)$, and second one via the use of additive factor $\tau(t)$. Both of these factors are affecting the effective growth rate of different factions in the generalized Lotka–Volterra system (GLV). We will now describe exogenous effects, that we will use in our GLV system.
@@
Regularization weights $\lambda_{\mathrm{axis}}=1$, $\lambda_{\mathrm{origin}}=1$ with a warmup of 100 epochs. In Appendix Fig. 5, we show the training results for the ABM-informed neural network on SIR dynamics. In this experiment (Appendix Fig. 5), we generate a single Erdős–Rényi contact network with node count $n=100$ and edge probability $p=0.05$. Ground-truth macroscopic curves $s(t)=\sum_{i}S_{i}(t)$, $i(t)=\sum_{i}I_{i}(t)$, and $r(t)=\sum_{i}R_{i}(t)$ are obtained by simulating an SIR model with infection and recovery rates $\beta=0.4$ and $\gamma=0.2$, respectively, using RK4 time stepping with $\mathrm{d}t=0.1$ up to $t_{\max}=60$. Initial conditions are set by infecting up to 50 nodes sampled from the largest connected component, with the remainder susceptible and no recoveries: $S_{i}(0)\in\{0,1\}$, $I_{i}(0)\in\{0,1\}$, $R_{i}(0)=0$, and $S_{i}(0)+I_{i}(0)=1$. Training uses a differentiable RK4 rollout in PyTorch on the fixed graph and minimizes an integral L1 path loss between the stacked state trajectories with a simple curriculum on the rollout horizon: starting at $t=1$ and increasing by $\Delta t=1$ every 50 epochs up to a cap (default $t_{\text{train}}\leq 30$). We add a regularization functional $\Omega(\phi_{1},\phi_{2})$ and optimize the sum of the data loss and $\Omega$. Appendix Fig. 5 also shows the training losses and regularizers at different epochs. We observe that the model learns the dynamics of the system, and the regularization terms enforce the desired properties.
After training, we simulate the learned ABM-informed neural network over the full horizon to compare $\tilde{s}(t),\tilde{i}(t),\tilde{r}(t)$ with the ground truth, inspect 1D slices of $\phi_{1}(0.5,x)$. Importantly, the structure of the SIR equations preserves node-wise mass: for each $i$, $\dot{S}_{i}+\dot{I}_{i}+\dot{R}_{i}=0$. With simplex-consistent initial conditions $S_{i}(0)+I_{i}(0)+R_{i}(0)=1$, the conservation law implies $S_{i}(t)+I_{i}(t)+R_{i}(t)=1$ for all $t$. Our architectural choices support this invariance, and the numerical scheme maintains it during simulation. In Hamiltonian neural networks, the neural parameterization enforces a symplectic vector field, which preserves a first integral (the Hamiltonian/energy) by construction. Analogously, in our epidemic setting the compartmental structure induces a conservation law (mass on each node). By designing the neural right-hand side so that $\dot{S}_{i}+\dot{I}_{i}+\dot{R}_{i}=0$, we add inductive bias that is the mass-conserving counterpart of the energy-preserving bias in HNNs, ensuring that neural dynamics remain on the probability simplex for each agent.
@@
All models are trained for 1000 epochs using the Adam optimizer with a learning rate of $10^{-4}$, minimizing the macroscopic $L_{1}$ loss derived from a differentiable RK4 rollout. ABM-NN Architecture: The ABM-NN decomposes dynamics into pairwise interactions ($\phi_{1}$) and self-recovery ($\phi_{2}$), both implemented as 3-layer MLPs (hidden dim: 32) with LeakyReLU activations. Baseline Architectures: The GCN baseline employs a stack of two GCNConv layers (hidden dimension 64) followed by a linear head. Similarly, GraphSAGE uses two SAGEConv layers (mean aggregation, hidden dimension 64) with a linear head. The GraphGPS transformer processes inputs through two GPS layers—each combining a local GCNConv with a 4-head global attention mechanism (hidden dimension 64)—before a final projection head.
@@
$\dot{\textbf{x}}=\underbrace{\phi_{x}(x_{i},y_{i},e_{i})}_{\text{self-interaction}}+\underbrace{\psi(\textbf{x},\textbf{A},e_{i},e_{j})}_{\text{interaction term}},$ (30)
$\dot{\textbf{y}}=\underbrace{\phi_{y}(y_{i},x_{i},e_{i})}_{\text{self-interaction}},$ (31)
@@
$\psi(x_{i},\textbf{x},\textbf{A},e_{i},e_{j})=\sum_{j}A_{ij}\,\phi_{2}(x_{i},x_{j},e_{i},e_{j}),$ (32)
where $\phi_{2}$ is a feed-forward neural network, and A is the adjacency matrix of the graph that is learned from the data.
@@
The shared self map $\phi_{x}$ follows the decomposition of Eq. (33),
@@
where $\phi_{g}(\cdot)$ is a feed-forward neural network that models annualized growth and $\phi_{d}(\cdot)$ is a feed-forward neural network that captures structural drifts. The latent macro network $\phi_{3}$ is a feed-forward neural network that models the dynamics of the macro channels. For economic modeling and macroeconomic data scarcity, we would like to have parsimonious models, we will use a shared feed-forward neural network for both $\phi_{x}(\cdot)$ and $\phi_{y}(\cdot)$ and $\psi$. Shared models correspond to the idea that economic laws follow some hidden economic universality laws or principles, and the same economic laws apply to all countries. However, in reality, each country has additional hidden variables or factors that are not captured by the shared model, and thus we will use FiLM-style adapters to model the country-specific conditional dynamics. Feature-wise Linear Modulation adapters (FiLM) were originally proposed for visual reasoning [perez2018film], and they have been shown to be effective in learning conditional models. In our case, we will use FiLM-style adapters to model the country-specific conditional macro dynamics.